# Multiplexed quantum repeaters with hot multimode alkali-noble gas memories

### v1

@@ Abstract
A solution to overcome fiber losses and achieve even longer distances is to use satellite-borne photon sources Yin et al. (2017a) or quantum memories Gündoğan et al. (2021), since photon absorption and decoherence in empty space are negligible. An entangled photon pair has been distributed to locations separated by $1200\ \mathrm{km}$ in a space-to-ground two-downlink channel Yin et al. (2017b), with losses as low as $64\ \mathrm{dB}$. However, satellite links are plagued by intermittency due to weather and satellite-orbit shadow time, which considerably reduce entanglement distribution rates. Hence, very long-lived quantum memories such as we propose here can alleviate these issues, storing entangled pairs when the satellite link is available Simon (2017), further increasing rates due to the temporal multiplexing enabled by the AFC. In this way, our proposed memory scheme could be incorporated in hybrid or satellite-based repeater schemes Gündoğan et al. (2021), thereby contributing towards the longstanding goal of building a quantum network spanning truly global distances.
@@
Acknowledgements.— The authors acknowledge Fundação para a Ciência e a Tecnologia (FCT-Portugal) through national funds and when applicable co-funding by EU funds under the Projects No. UIDB/50008/2020, PTDC/FIS-OUT/3882/2020, and Contract Nos. CEECIND/00401/2018 and CEECIND/CP1653/CT0002.

### v2

@@ Abstract
A solution to overcome fiber losses and achieve even longer distances is to use satellite-borne photon sources [54] or quantum memories [55], since photon absorption and decoherence in empty space are negligible. An entangled photon pair has been distributed to locations separated by $1200\ \mathrm{km}$ in a space-to-ground two-downlink channel [56], with losses as low as $64\ \mathrm{dB}$. However, satellite links are plagued by intermittency due to weather and satellite-orbit shadow time, which considerably reduce entanglement distribution rates. Hence, very long-lived quantum memories such as we propose here can alleviate these issues, storing entangled pairs when the satellite link is available [53], further increasing rates due to the temporal multiplexing enabled by the AFC. In this way, our proposed memory scheme could be incorporated in hybrid or satellite-based repeater schemes [55], thereby contributing towards the longstanding goal of building a quantum network spanning truly global distances.
Code Availability. The code that supports the findings of this Letter is openly available [57].
@@
Acknowledgements. The authors acknowledge Fundação para a Ciência e a Tecnologia (FCT-Portugal) through national funds and, when applicable, co-funding by EU funds under the Projects No. UIDB/50008/2020, PTDC/FIS-OUT/3882/2020, and Contract Nos. CEECIND/00401/2018 and CEECIND/CP1653/CT0002.